# Text-Guided Layer Fusion Mitigates Hallucination in Multimodal LLMs

### v1

@@ Abstract
Multimodal large language models (MLLMs) typically rely on a single late-layer feature from a frozen vision encoder, leaving the encoder’s rich hierarchy of visual cues under-utilized. MLLMs still suffer from visually ungrounded hallucinations, often relying on language priors rather than image evidence. While many prior mitigation strategies operate on the text side, they leave the visual representation unchanged and do not exploit the rich hierarchy of features encoded across vision layers. Existing multi-layer fusion methods partially address this limitation but remain static, applying the same layer mixture regardless of the query. In this work, we introduce TGIF (Text-Guided Inter-layer Fusion), a lightweight module that treats encoder layers as depth-wise “experts” and predicts a prompt-dependent fusion of visual features. TGIF follows the principle of direct external fusion, requires no vision-encoder updates, and adds minimal overhead. Integrated into LLaVA-1.5-7B, TGIF provides consistent improvements across hallucination, OCR, and VQA benchmarks, while preserving or improving performance on ScienceQA, GQA, and MMBench. These results suggest that query-conditioned, hierarchy-aware fusion is an effective way to strengthen visual grounding and reduce hallucination in modern MLLMs. Our code will be available at: https://github.com/Linchenchen/TGIF. Refer to caption Figure 1: Comparison of layer fusion designs in MLLMs. (a) MLP Connector: Uses only the penultimate layer of the vision encoder, mapping global visual tokens through a simple projection. (b) Dense Connector: Aggregates multi-layer visual features via concatenation or downsampling before projection, enriching semantics but with a fixed fusion pattern. (c) MMFuser: Retrieves shallow-layer features using deep-layer queries (Q–K/V attention) to capture local details. (d) Proposed TGIF: Introduces a text-guided layer router that dynamically reweights features from multiple layers based on the input query, enabling adaptive, context-aware visual fusion.
@@ 1 Introduction
Multimodal large language models (MLLMs) have recently achieved impressive progress on visual question answering, captioning, and open-ended multimodal dialogue by combining the reasoning ability of large language models (LLMs) with the perceptual capacity of pretrained vision encoders [llava, li2023blip, instructblip]. Most state-of-the-art systems adopt a modular design: a frozen vision encoder (e.g., CLIP ViT), a lightweight connector, and a powerful LLM decoder. The connector projects visual embeddings into the LLM’s token space, enabling the model to jointly process image and text tokens for downstream reasoning.
Despite these advances, modern MLLMs still frequently produce confident but visually ungrounded descriptions, a phenomenon broadly referred to as hallucination [rohrbach2019objecthallucinationimagecaptioning, liu2024surveyhallucinationlargevisionlanguage]. In the vision–language setting, hallucination typically manifests as objects, attributes, or relations that are plausible under language priors but inconsistent with the input image. This issue is especially severe for detail-oriented tasks (e.g., OCR, small object recognition), where high-level semantic features alone are insufficient to support fine-grained grounding.
Prior work tackles hallucination from two main angles. Training-based approaches improve alignment via additional instruction tuning, contrastive fine-tuning, or RLHF [gunjal2024detectingpreventinghallucinationslarge, yu2024rlhfvtrustworthymllmsbehavior], but they are often data- and compute-intensive. Training-free approaches instead modify decoding or inference without retraining the model. Methods such as VCD [leng2023mitigatingobjecthallucinationslarge], VTI [liu2024reducinghallucinationsvisionlanguagemodels], OPERA [huang2024opera], FarSight [tang2025seeingfarclearlymitigating], and PerturboLLaVA [chen2025perturbollavareducingmultimodalhallucinations] mitigate hallucination via contrastive calibration, latent intervention, causal masking, or over-trust penalties. However, these strategies primarily operate on the text side: they adjust the decoder or filter responses post hoc, while the underlying visual representation typically remains a single, fixed layer of the vision encoder passed through an MLP projector.
Meanwhile, transformer-based vision encoders such as CLIP are known to build a rich hierarchy of visual abstractions across layers: shallow layers preserve edges, textures, and local geometry, whereas deeper layers encode high-level semantics aligned with text [clip, gandelsman2024interpretingclipsimagerepresentation]. Recent works have begun to exploit this hierarchy by fusing features from multiple depths. DenseConnector [yao2024denseconnectormllms] concatenates or downsamples features from selected layers, and MMFuser [cao2024mmfusermultimodalmultilayerfeature] retrieves shallow-layer details using deep features as queries. A recent systematic study further shows that direct, external fusion of multi-layer visual features at the input stage yields the most stable performance across architectures [lin2025multilayervisualfeaturefusion]. However, existing fusion schemes are static: the mixture of layers is fixed once chosen, independent of the question, and tends to overemphasize globally aligned semantics even when the query requires local evidence.
In this work, we argue that hallucination in MLLMs is fundamentally linked to how visual features are selected and exposed to the LLM. We propose TGIF (Text-Guided Inter-layer Fusion), a dynamic routing framework that treats the layers of a frozen vision encoder as a pool of specialized “experts” and adaptively fuses them based on the input query. Concretely, TGIF introduces a lightweight router inside the multimodal projector that takes text (and optionally a global image feature) as input and outputs a soft distribution over all ViT layers. These learned weights are then used to form a query-conditioned fused visual representation, which is projected into the LLM token space using a standard MLP connector (Fig. 2). Following the principle of direct external fusion [lin2025multilayervisualfeaturefusion], TGIF keeps the vision encoder frozen and preserves the token budget, while allowing the layer mixture to vary per prompt.
We instantiate TGIF on top of LLaVA-1.5 [li2023llava] with CLIP-ViT-L/14 and Vicuna-7B, and evaluate it across three benchmark families: hallucination, OCR, and general VQA. On hallucination-focused evaluations, TGIF achieves state-of-the-art performance among 7B-scale LLaVA variants, improving POPE accuracy from 86.85% to 87.91% and boosting HallusionBench All Accuracy from 46.90% to 49.94%, outperforming both decoding-based baselines (VCD, OPERA, VTI, FarSight, PerturboLLaVA) and larger 13B models. On OCRBench, TGIF improves the final score from 297 to 313 (+16), and provides consistent gains on TextVQA. At the same time, it maintains or improves overall performance on ScienceQA, GQA, and MMBench, indicating that better grounding does not come at the expense of high-level reasoning.
Our main contributions are three-fold:
@@
  • We identify a key limitation of current multimodal LLMs: visual tokens are typically drawn from a single, late-layer representation, which is poorly suited for detail-sensitive grounding and exacerbates hallucination under strong language priors.
@@
  • We propose TGIF, a text-guided inter-layer fusion module that dynamically reweights CLIP layers per query, following the direct external fusion principle while remaining parameter- and token-efficient. We explore both text-only and multimodal routers, and introduce an entropy-based load-balancing loss to prevent expert collapse.
@@
  • We demonstrate that TGIF substantially improves hallucination robustness and fine-grained visual perception on POPE, HallusionBench, OCRBench, and TextVQA, while preserving competitive performance on general reasoning benchmarks. Qualitative analysis of router behavior further shows that TGIF learns semantically meaningful depth-selection patterns across task types.
@@ 2.1 Multimodal Large Language Models
Multimodal large language models (MLLMs) integrate the reasoning ability of LLMs with the perceptual capacity of pretrained vision encoders. Most follow a modular design with a frozen vision encoder, a lightweight connector, and a powerful LLM decoder. The connector projects visual embeddings into the text space, enabling cross-modal alignment essential for grounded reasoning. Early designs used simple MLP projectors [llava], while later approaches like BLIP-2 [li2023blip] and InstructBLIP [instructblip] employ query-based modules for salient visual token extraction.
2.2 Hallucination in MLLMs
MLLMs frequently generate confident yet ungrounded content—a phenomenon known as hallucination [rohrbach2019objecthallucinationimagecaptioning, liu2024surveyhallucinationlargevisionlanguage]. Training-based mitigation strategies rely on additional instruction tuning, contrastive fine-tuning, or RLHF [gunjal2024detectingpreventinghallucinationslarge, yu2024rlhfvtrustworthymllmsbehavior], but they are compute- and data-intensive. Training-free methods instead modify decoding or inference. Recent works such as VCD [leng2023mitigatingobjecthallucinationslarge], VTI [liu2024reducinghallucinationsvisionlanguagemodels], OPERA [huang2024opera], FarSight [tang2025seeingfarclearlymitigating], and PerturboLLaVA [chen2025perturbollavareducingmultimodalhallucinations] mitigate hallucination via contrastive calibration, token intervention, or causal masking. However, these approaches primarily act on the text side, leaving the underlying vision-language alignment unchanged. Our work complements them by targeting hallucination at the feature level through text-guided multi-layer fusion.
@@ 2.3 Multi-Layer Visual Feature Fusion
Transformer-based vision encoders like CLIP exhibit a hierarchical structure where deeper layers capture semantic abstraction and intermediate layers preserve fine-grained spatial cues [clip, gandelsman2024interpretingclipsimagerepresentation]. DenseConnector [yao2024denseconnectormllms] concatenates features from multiple layers, while MMFuser [cao2024mmfusermultimodalmultilayerfeature] retrieves shallow-layer details using deep-layer queries. These static strategies enrich representations but cannot adapt fusion to each query. Recent work also systematically analyzes layer integration strategies, showing that direct external fusion yields the most stable performance [lin2025multilayervisualfeaturefusion]. Building on these insights, TGIF performs text-guided inter-layer fusion, dynamically reweighting visual features according to the input query to improve grounding and reduce hallucination.
@@ 3 Method
Our approach is situated within a standard Vision Language Model (VLM) architecture, which comprises a vision encoder (e.g., CLIP ViT [clip]), a language model (e.g., Vicuna [vicuna]), and a multimodal projector that maps visual features into the language model’s embedding space. Our proposed text-guided layer selection module is designed as a core component of this multimodal projector. It takes as input the full stack of hidden states from the all vision encoder layers, along with features derived from the text prompt, to produce a dynamically tailored visual representation for the LLM.
@@ 3.1 Text-Guided Layer Selection
As established in prior work [chen2025multimodallanguagemodelsbetter], different CLIP layers capture distinct semantic information, with shallow layers encoding textures and spatial details while deeper layers align more with global semantics. This suggests that the layers can be viewed as a pool of specialized “experts”. Borrowing concepts from the Mixture-of-Experts (MoE) paradigm, we treat each layer as an expert and propose a dynamic layer selection framework, which we name TGIF (Text-Guided Inter-layer Fusion). In this framework, a text-guided “router” learns to generate weights to select and fuse the most relevant layer experts for a given task. We explore two architectures for the TGIF router, one with text-only input and other with multimodal input. The model architecture and framework is shown in Fig 2. This design allows adaptive, query-conditioned fusion of depth-wise visual cues, improving both grounding and fine-grained detail understanding.
@@ 3.1.1 Text-Guided MLP Router
Our baseline approach uses a lightweight MLP-based router to predict layer importance scores based solely on the textual prompt. This router learns a direct mapping from the question’s semantics to the relevance of different vision layers.
Let $\mathbf{f}_{\text{text}}\in\mathbb{R}^{D_{t}}$ denote the pooled text embedding from the LLM. Let $\{\mathbf{F}_{l}\in\mathbb{R}^{P\times D_{v}}\}_{l=1}^{L}$ be the set of patch-level visual features from all $L$ layers of the vision encoder. The MLP selector first predicts unnormalized logits for each layer:
@@ 3.1.2 Multimodal MLP Router
where $\mathbf{W}_{t}\in\mathbb{R}^{D_{t}\times D_{p}}$ and $\mathbf{W}_{v}\in\mathbb{R}^{D_{v}\times D_{p}}$ are learnable projection matrices. This combined multimodal feature vector is then used by the MLP to predict the layer weights, following the same process as in Equations 1-3.
@@ 3.2 Load Balancing Loss
A common challenge when training MoE-style routers is the tendency for the router to converge to a state where it consistently selects the same few “safe” experts (in our case, layers), leading to “expert starvation” [shazeer2017outrageouslylargeneuralnetworks]. To mitigate this and encourage the router to utilize a more diverse set of layers, we incorporate a modified auxiliary load balancing loss into our total training objective.
@@
Because the nature of textual prompts differs between pretraining and instruction tuning dataset, we apply different $\lambda$ values across stages. During pretraining, the router often receives generic prompts (e.g., “Describe the image”), which provide limited textual guidance. We thus strengthen the visual signal by applying a slightly larger $\lambda$ to encourage exploration of multiple layers. During fine-tuning, prompts are task-oriented and semantically rich (e.g., “What number is written on the sign?”). Here, we reduce $\lambda$ to allow the router to focus on discriminative, text-conditioned layer selection.
@@ 4 Experiments
POPE HallusionBench TextVQA OCRBench ScienceQA GQA MMBench
@@
+ TGIF (Text-Only MLP) 87.30 49.95 58.93 31.50 68.07 62.48 65.97
+ TGIF (Multimodal MLP) 86.26 57.31 59.09 29.90 69.72 62.37 65.46
+ TGIF ($\lambda{=}0.01$ pretrain-only) 87.91 48.68 58.98 31.30 70.10 62.58 66.40
Table 1: Comparison across VLM benchmarks. Hallucination Benchmarks: POPE and HallusionBench evaluate factual grounding and object hallucination. OCR Benchmarks: TextVQA and OCRBench assess text recognition and fine-grained perception. Overall Performance: ScienceQA, GQA, and MMBench measure general reasoning and instruction-following. Best per metric in bold; second best underlined.
Method Accuracy $\uparrow$ F1 Score $\uparrow$
@@
Method Params All Acc. $\uparrow$
@@
4.1 Experimental Setting
4.1.1 Implementation Details
We implement our proposed framework, TGIF, on top of the publicly available LLaVA-1.5 [li2023llava] codebase. To ensure a fair comparison, our primary experiments maintain consistency with LLaVA-1.5 by employing CLIP-ViT-L/14-336px [clip] as the vision encoder and Vicuna-7B [vicuna] as the LLM.
4.1.2 Training Recipe
We train all models on 8 NVIDIA H100-80G GPUs following the two-stage training paradigm of LLaVA-1.5 [li2023llava].
Stage 1 (Feature Alignment Pretraining). In this stage, both the vision encoder and the LLM remain frozen. We train only the TGIF components, the layer router and MLE connector module, on a filtered subset of 558K image–text pairs from CC3M. This stage aligns multi-layer visual representations with the LLM’s embedding space and allows the router to learn an initial prompt-adaptive layer selection policy. We use a learning rate of 1e-3 and a global batch size of 256.
Stage 2 (Instruction Finetuning). Following pretraining, we fine-tune the LLM together with the TGIF projector while keeping the vision encoder frozen. Here the dataset we use is the 665K multi-turn conversations for instruction fine-tuning. This phase refines the model’s conversational and reasoning abilities using dynamically fused visual features. The learning rate is set to 2e-5 with a batch size of 128.
We also apply stage-specific load balancing coefficients (Sec. 3.2) to encourage diverse layer usage during pretraining and more discriminative text-conditioned routing during fine-tuning.
4.1.3 Evaluation Benchmarks
To comprehensively evaluate our TGIF framework, we benchmark its performance across a diverse set of multimodal tasks spanning hallucination detection, fine-grained OCR reasoning, and general visual question answering. All evaluations are conducted using the standardized VLMEvalKit [duan2024vlmevalkit] platform to ensure consistency and comparability across models.
Hallucination Benchmarks. To assess grounding faithfulness, we adopt two representative hallucination benchmarks: HallusionBench (HB) [hallusionbench] and POPE [Li-hallucination-2023]. HallusionBench probes visual factuality by testing a model’s ability to reject implausible object claims, while POPE reformulates hallucination detection as a binary classification task, quantifying the model’s awareness of object existence.
OCR Benchmarks. To evaluate text recognition and detail-sensitive reasoning, we include TextVQA [singh2019vqamodelsread] and the comprehensive OCRBench [Liu_2024], which covers scene-text, document-text, and key information extraction subtasks. These benchmarks reveal the model’s ability to retrieve and interpret fine-grained textual cues—an area where hallucination often manifests due to overreliance on semantic priors.
General Reasoning and Overall Benchmarks. For overall multimodal reasoning performance, we report results on widely used visual QA and instruction-following datasets, including MMBench (MMB) [liu2024mmbench], ScienceQA [lu2022learn] and GQA [ainslie2023gqatraininggeneralizedmultiquery]. Together, these benchmarks measure both the factual grounding and generalization capability of TGIF across visual domains and task types.
4.1.4 Baseline Methods
We evaluate TGIF against two complementary categories of baselines: (1) hallucination-mitigation methods that focus on decoding and inference, and (2) multi-layer fusion architectures that enhance visual representations.
Hallucination Mitigation Methods. We compare TGIF with five representative training-free approaches that operate on the decoding side. VCD [leng2023mitigatingobjecthallucinationslarge] introduces visual contrastive decoding, contrasting outputs from original and distorted inputs to reduce unimodal bias. VTI [liu2024reducinghallucinationsvisionlanguagemodels] stabilizes cross-modal interactions by injecting visual and textual interventions in latent space during inference. OPERA [huang2024opera] applies an over-trust penalty and retrospection-allocation decoding to discourage over-attention to summary tokens. FarSight [tang2025seeingfarclearlymitigating] improves token propagation through causal masking, mitigating attention drift toward outlier tokens. PerturboLLaVA [chen2025perturbollavareducingmultimodalhallucinations] enhances robustness by perturbing the visual embedding space during training to counteract language priors. These methods primarily target textual hallucinations during generation and provide a strong benchmark for comparison on HallusionBench and POPE.
Layer Fusion Baselines. For fair evaluation of visual representation improvements, we also compare with Dense Connector [yao2024denseconnectormllms] and MMFuser [cao2024mmfusermultimodalmultilayerfeature], which aggregate visual information from multiple layers of the vision encoder. Dense Connector concatenates or downsamples features from selected depths, while MMFuser retrieves shallow-layer details using deep features as queries. These architectures serve as strong baselines for evaluating how TGIF’s dynamic, text-guided fusion enhances multimodal grounding and general reasoning.
@@ 4.1.4 Baseline Methods
LLaVA1.5-7B 160 117 15 5 0 297
TGIF (ours) 162 121 24 6 0 313
Table 4: OCRBench Comparison. Recog.:Text Recognition; $VQA^{S}$:Scene Text-centric VQA; $VQA^{D}$:Document-oriented VQA; KIE: Key Information Extraction; HMER:Handwritten Math Expression Recognition. TGIF improves over LLaVA-1.5-7B by +16 on the Final score.
4.2 Experimental Results
@@
We further analyze TGIF’s performance on hallucination mitigation, fine-grained perception, and general reasoning to understand its behavior across tasks. For detailed performance comparisons, we adopt the best-performing configuration, TGIF with a pretraining-only load-balancing coefficient of $\lambda=0.01$.
4.2.1 Hallucination mitigation

### v2

@@ Abstract
Multimodal large language models (MLLMs) often generate confident but visually ungrounded responses, a phenomenon known as hallucination. While prior mitigation methods primarily operate at decoding time or rely on additional training, they typically expose the language model to a fixed, late-layer visual representation, overlooking the rich hierarchy encoded by vision transformers. We show that hallucination behavior is strongly influenced by the depth of visual features provided to the LLM, and that no single layer is optimal across queries. To address this, we propose Text-Guided Inter-layer Fusion (TGIF), a lightweight architectural module that dynamically reweights visual features across transformer layers based on the input text, without modifying the vision encoder or increasing the token budget. Experiments on hallucination, OCR, and general VQA benchmarks demonstrate that TGIF substantially improves visual grounding and hallucination robustness while preserving strong overall reasoning performance.
@@ 1  Introduction
Multimodal large language models (MLLMs) combine the reasoning capabilities of large language models (LLMs) with pretrained vision encoders, achieving impressive results in diverse visual tasks [26, 19, 9]. However, MLLMs face persistent hallucination challenge, generating confident but ungrounded content that is plausible under language priors yet inconsistent with the image [24]. This is particularly pronounced in detail-oriented tasks where high-level semantic features alone are insufficient for precise grounding. Most existing hallucination mitigation methods intervene after training or at inference time [13, 46, 16, 27, 14, 36, 5]. While effective, these approaches primarily operate on the text generation process and leave the underlying visual representations unchanged.
Figure 1: Impact of ViT Layer Injection on Hallucination Behavior. The figure illustrates three scenarios: (1) Shallow layer injection leads to a lack of semantic understanding. (2) Middle layer injection captures sufficient object-level features for an accurate answer. (3) Deep layer injection can lead to hallucination or misinterpretation of the global context.
A key but underexplored factor is the depth of visual representations exposed to the language model. Vision transformers such as CLIP encode a hierarchy of visual abstractions across layers, ranging from low-level spatial cues to high-level semantic concepts. However, most MLLMs expose only a single, fixed (typically late) vision layer to the LLM via an MLP projector. Figure 1 presents a representative example in which the baseline LLaVA model hallucinates, demonstrating that injecting visual features from different depths, with all other components held fixed, results in qualitatively different behaviors ranging from under-recognition to hallucination.
Recent works begin to highlight the role of visual encoder layers in MLLMs, suggesting that shallow and middle layers exhibit unique strengths over deep layers in certain sub-tasks [6]. VaLiD [27] and LayerCD [5] collectively suggest that hallucination behavior is sensitive to the choice of visual feature depth. To the best of our knowledge, however, no prior work has addressed this issue by explicitly modifying the multimodal architecture to dynamically control which visual representations are exposed to the LLM.
Motivated by this observation, we hypothesize that hallucination in MLLMs is strongly influenced by how visual features are selected and exposed to the language model, and that adapting this selection based on the input prompt can improve visual grounding. We propose TGIF (Text-Guided Inter-layer Fusion), a lightweight architectural module that dynamically reweights visual features across all layers of a frozen vision encoder according to the textual query (Fig. 2). This design preserves the pretrained backbone and token budget while enabling prompt-dependent control over visual abstraction.
We instantiate TGIF on top of LLaVA-1.5 and evaluate it on hallucination, OCR, and general VQA benchmarks. TGIF consistently improves hallucination robustness and fine-grained visual perception while maintaining competitive performance on general reasoning tasks, demonstrating that better visual grounding can be achieved through dynamic representation-level control rather than post hoc decoding intervention. Our main contributions are three-fold:
@@
  • We identify a limitation of current MLLMs: visual tokens are typically drawn from a single late-layer representation, which is poorly suited for detail-sensitive grounding and exacerbates hallucination under strong language priors.
@@
  • We propose TGIF, a text-guided inter-layer fusion module that dynamically reweights CLIP layers per query while remaining parameter- and token-efficient. We explore both text-only and multimodal routers and introduce an entropy-based load-balancing loss.
@@
  • We show that TGIF improves hallucination robustness and fine-grained visual perception on POPE, HallusionBench, OCRBench, and TextVQA, while maintaining competitive performance on general reasoning benchmarks.
@@ 2.1  Multimodal Large Language Models
Multimodal large language models (MLLMs) integrate the reasoning ability of LLMs with the perceptual capacity of pretrained vision encoders. Most follow a modular design with a frozen vision encoder, a lightweight connector, and a powerful LLM decoder. The connector projects visual embeddings into the text space, enabling cross-modal alignment essential for grounded reasoning. Early designs used simple MLP projectors [26], while later approaches like BLIP-2 [19] and InstructBLIP [9] employ query-based modules for salient visual token extraction. Despite architectural differences, a common assumption across these models is that penultimate visual encoder layer is sufficient for all queries.
2.2  Hallucination Mitigation in MLLMs
Multimodal large language models (MLLMs) are prone to generating confident yet visually ungrounded responses, commonly referred to as hallucinations [32, 24]. Training-based mitigation approaches typically rely on additional instruction tuning, contrastive fine-tuning, or reinforcement learning from human feedback (RLHF) [13, 46], but introduce substantial computational and data costs. Training-free methods primarily operate at the decoding stage, for example by calibrating token probabilities, modifying attention patterns, or applying causal masking [16, 27, 14, 36, 5]. While effective in practice, these approaches largely act on the language generation process, without explicitly addressing how visual representations are selected and exposed to the language model.
Recent studies begin to highlight the role of visual representations in hallucination. VaLiD [41] attributes hallucination to distortions in visual encoding and mitigates it via visual layer fusion combined with contrastive decoding guided by uncertainty. LayerCD [39] further demonstrates that hallucinated outputs can be suppressed by contrasting decoding behaviors induced by shallow versus deep visual features. Despite differing hypotheses, these works collectively suggest that hallucination behavior is sensitive to the choice of visual feature depth.
@@ 2.3  Multi-Layer Visual Feature Fusion
Transformer-based vision encoders like CLIP exhibit a hierarchical structure where deeper layers capture semantic abstraction and intermediate layers preserve fine-grained spatial cues [31, 11]. DenseConnector [42] concatenates features from multiple layers, while MMFuser [4] retrieves shallow-layer details using deep-layer queries. These static strategies enrich representations but cannot adapt fusion to each query. Recent work also systematically analyzes layer integration strategies, showing that direct external fusion yields the most stable performance [21]. Building on these insights, TGIF performs text-guided inter-layer fusion, dynamically reweighting visual features according to the input query to improve grounding and reduce hallucination. Figure 6 in the Appendix provides a schematic comparison between static multi-layer fusion designs and our proposed text-guided routing.
@@ 3  Method
Our approach is situated within a standard Vision Language Model (VLM) architecture, which comprises a vision encoder (e.g., CLIP ViT [31]), a language model (e.g., Vicuna [7]), and a multimodal projector that maps visual features into the language model’s embedding space. Our proposed text-guided layer selection module is designed as a core component of this multimodal projector. It takes as input the full stack of hidden states from the all vision encoder layers, along with features derived from the text prompt, to produce a dynamically tailored visual representation for the LLM.
@@ 3.1  Text-Guided Layer Selection
As established in prior work [6], different CLIP layers capture distinct semantic information, with shallow layers encoding textures and spatial details while deeper layers align more with global semantics. This suggests that the layers can be viewed as a pool of specialized “experts”. Inspired by the Mixture-of-Experts (MoE) paradigm, we treat each layer as an expert and propose a dynamic layer selection framework, called TGIF (Text-Guided Inter-layer Fusion). In this framework, a text-guided “router” learns to generate weights to select and fuse the most relevant layer experts for a given task. We explore two architectures for the TGIF router, one with text-only input and other with multimodal input. The model architecture and framework is shown in Fig 2. This design allows adaptive, query-conditioned fusion of depth-wise visual cues, improving both grounding and fine-grained detail understanding.
@@ 3.1.1  Text-Guided MLP Router
Our baseline approach uses a lightweight MLP-based router to predict layer importance scores based solely on the textual prompt. This router learns a direct mapping from the question’s semantics to the relevance of different vision layers. Let $\mathbf{f}_{\text{text}}\in\mathbb{R}^{D_{t}}$ be the pooled text embedding from the LLM, $\{\mathbf{F}_{l}\in\mathbb{R}^{P\times D_{v}}\}_{l=1}^{L}$ be the set of patch-level visual features from all $L$ layers of the vision encoder. The MLP selector first predicts unnormalized logits for each layer:
@@ 3.1.2  Multimodal MLP Router
where $\mathbf{W}_{t}\in\mathbb{R}^{D_{t}\times D_{p}}$ and $\mathbf{W}_{v}\in\mathbb{R}^{D_{v}\times D_{p}}$ are learnable projection matrices. This combined multimodal feature vector is then used by the MLP to predict the layer weights, following the same process as in Equations 1-3. A schematic comparison between the text-only and multimodal routers is provided in Appendix B (Fig. 7).
@@ 3.2  Load Balancing Loss
A common challenge when training MoE-style routers is the tendency for the router to converge to a state where it consistently selects the same few “safe” experts (in our case, layers), leading to “expert starvation” [33]. To mitigate this and encourage the router to utilize a more diverse set of layers, we incorporate a modified auxiliary load balancing loss into our total training objective.
@@
Because the nature of textual prompts differs between pretraining and instruction tuning dataset, we apply different $\lambda$ values across stages. During pretraining, the router often receives generic prompts (e.g., “Describe the image”), which provide limited textual guidance. We thus strengthen the visual signal by applying a slightly larger $\lambda$ to encourage exploration of multiple layers. During fine-tuning, prompts are task-oriented and semantically rich (e.g., “What number is written on the sign?”). So we reduce $\lambda$ to allow the router to focus on discriminative, text-conditioned layer selection.
@@ 4  Experiments
4.1  Experimental Setting
4.1.1  Implementation Details
We implement our proposed framework, TGIF, on top of the publicly available LLaVA-1.5 [17] codebase. To ensure a fair comparison, our primary experiments maintain consistency with LLaVA-1.5 by employing CLIP-ViT-L/14-336px [31] as the vision encoder and Vicuna-7B [7] as the LLM.
4.1.2  Training Recipe
All models are trained on 8 NVIDIA H100-80G GPUs following the two-stage training paradigm of LLaVA-1.5 [17], consisting of Stage 1: Feature Alignment Pretraining and Stage 2: Instruction Finetuning. For a fair comparison, we adopt the same training datasets and data splits as LLaVA-1.5. Implementation details for each stage are provided in Sec. C. We also apply stage-specific load balancing coefficients (Sec. 3.2) to encourage diverse layer usage during pretraining and more discriminative text-conditioned routing during fine-tuning.
4.1.3  Evaluation Benchmarks
To comprehensively evaluate our TGIF framework, we benchmark its performance across a diverse set of multimodal tasks spanning hallucination detection, fine-grained OCR reasoning, and general visual question answering. All evaluations are conducted using the standardized VLMEvalKit [10] platform to ensure consistency and comparability across models.
Hallucination Benchmarks. To assess grounding faithfulness, we adopt two representative hallucination benchmarks: HallusionBench (HB) [22] and POPE [20]. HallusionBench probes visual factuality by testing a model’s ability to reject implausible object claims, while POPE reformulates hallucination detection as a binary classification task, quantifying the model’s awareness of object existence.
OCR Benchmarks. To evaluate text recognition and detail-sensitive reasoning, we include TextVQA [35] and the comprehensive OCRBench [29], which covers scene-text, document-text, and key information extraction subtasks. These benchmarks reveal the model’s ability to retrieve and interpret fine-grained textual cues—an area where hallucination often manifests due to overreliance on semantic priors.
General Reasoning and Overall Benchmarks. For overall multimodal reasoning performance, we report results on widely used visual QA and instruction-following datasets, including MMBench (MMB) [28], ScienceQA [30] and GQA [1]. Together, these benchmarks measure both the factual grounding and generalization capability of TGIF across visual domains and task types.
4.1.4  Baseline Methods
We evaluate TGIF against two complementary categories of baselines:
Decoding-based Hallucination Mitigation. VCD [16], VTI [27], OPERA [14], FarSight [36], and PerturboLLaVA [5] intervene at decoding or inference time without modifying visual representations.
Multi-layer Fusion Baselines. DenseConnector [42] and MMFuser [4] aggregate features from multiple vision encoder layers using static fusion strategies.
4.2  Main Results
Table 1: Comparison across VLM benchmarks. Hallucination: POPE and HallusionBench(HB). OCR: TextVQA and OCRBench. Overall: ScienceQA(SQA), GQA, MMBench. Best in bold, second best underlined. $\dagger$ Our final TGIF model; other TGIF variants explore router design choices.
@@ 4.2  Main Results
POPE HB TextVQA OCRBench SQA GQA MMBench
@@
+ TGIF (Text-only router) 87.30 49.95 58.93 31.50 68.07 62.48 65.97
+ TGIF (Multimodal router) 86.26 57.31 59.09 29.90 69.72 62.37 65.46
+ TGIF ($\lambda{=}0.01$) 87.91 48.68 58.98 31.30 70.10 62.58 66.40
Method Acc $\uparrow$ F1 $\uparrow$
@@
Table 3: POPE-Adversarial depth analysis. Static depth choices exhibit complementary failure modes, while TGIF achieves the best F1-score via text-guided inter-layer fusion.
[…]
LLaVA 0.851 0.899 0.792 0.852
+Early-only 0.804 0.963 0.632 0.763
+Mid-only 0.840 0.942 0.725 0.819
+Late-only 0.837 0.800 0.899 0.847
+Uniform 0.838 0.957 0.707 0.813
+TGIF(ours) 0.860 0.881 0.832 0.856
Method Params All Acc $\uparrow$
@@
Table 5: OCRBench comparison. Recog.:Text Recognition; $VQA^{S}$:Scene Text-centric VQA; $VQA^{D}$:Document-oriented VQA; KIE: Key Information Extraction; HMER:Handwritten Math Expression Recognition.
@@
LLaVA-1.5-7B 160 117 15 5 0 297
+TGIF (ours) 162 121 24 6 0 313
@@
We further analyze TGIF’s performance on hallucination mitigation, fine-grained perception, and general reasoning to understand its behavior across tasks. For detailed performance comparisons, we adopt the best-performing configuration, TGIF with a pretraining-only load-balancing coefficient of $\lambda=0.01$. Unless otherwise specified, TGIF refers to this final configuration ($\lambda{=}0.01$) throughout all tables and figures.
4.3  Hallucination Mitigation
4.3.1  POPE and HallusionBench